Find values for m and b in the following system so that the solution to the system
is (-3, 4).
5x + 7y = b
mx + y = 22

Answer by stanbon(75887)

Find values for m and b in the following system so that the solution to the system
is (-3, 4).
5x + 7y = b
mx + y = 22
-----------------
Substitute for x and y to get:
5*-3 + 7*4 = b
m*-3 + 4 = 22
------------------------
b = -15+28 = 13
-3m = 18
m = -6
-------------------
Solution: b = 13 ; m = -6
=================================
Cheers,
Stan H.